# Beyond Copenhagen: Following the Trail of Decoherence in Feynman’s Light Microscope

### v2

@@ XII Conclusion
Decoherence is the penalty we pay for trying to separate out one part of a non-separable entangled state. Wonderfully, this “illegal” separation accomplishes what Copenhagen clumsily attempted by imposing its shifty split. The quantum freedom to choose different bases creates an ambiguity as to the facts of any unfolding story. In Feynman’s light microscope, is the atom following a classical single-path trajectory or a non-classical two-path trajectory? With full environmental information, both narratives are legitimate and both make experimentally verifiable predictions. But without ever imposing a microscopic/macroscopic distinction, decoherence causes plain quantum mechanics to pick out the classical single-path storyline as the only one which can be recorded and used to make predictions in a complex environment. These decohered storylines each contain a set of unambiguous and causally connected facts.
@@
Alternatively, informational interpretations understand the wavefunction to be only a tool for making predictions. If the wavefunction never described microscopic reality, there is no basis for saying that any of the decohered branches predicted by unitary evolution actually exist, other than the one we are experiencing.
As a practical matter, MWI and informational approaches use identical equations, naturally include decoherence in identical ways, and make identical predictions for experiments. Opinions differ on whether the distinction is therefore an unimportant matter of philosophical preference [31, 37]. However, MWI and informational theories are together testable. For instance, if interference fringes of a big object were found to wash out for an inexplicable reason, then the underpinning unitary evolution would be falsified.
It is of course possible that quantum mechanics, and its extension quantum field theory, need to be modified in some fundamental way. Advocates point out that quantum theory has not yet been reconciled with gravity, and that there are still orders of magnitude between where it has been tested and the scale of macroscopic detectors. Alternate theories deserve to be tested when possible, but we should keep in mind that unitary evolution of Schrödinger’s equation has continued to prove successful under increasingly stringent conditions [2, 3, 4, 5]. Quantum mechanics has no crisis which modifications would address, other than our human crisis of being forced to question cherished notions of reality or science.
Half a century ago, physicists could believe that the wavefunction described the reality of small things and that there was only one macroscopic reality. But after the development of decoherence theory and experimental scrutiny at relatively large mass scales, we can no longer rely on wavefunction collapse to provide that comfortable scenario. We must either say that plain quantum mechanics correctly predicts detector clicks but makes no statements about microscopic reality, or that it predicts a vast number of macroscopic realities besides the one we are currently experiencing.
@@ Acknowledgements.
We gratefully acknowledge stimulating and clarifying discussions with Jerry Gabrielse, Tim Kovachy, Jay Lawrence, and Bill Wooters.
@@ Appendix A Calculation of Interference Patterns
We are interested in obtaining atom patterns for a detector closer than the far-field limit, such that the single-slit patterns are somewhat separated from one another. To accommodate this requirement, we use a path integral approach. The phase acquired by an atom momentum eigenstate traveling from a source position $x^{\prime}$ to a detector position $x$ given by
$\Phi_{0}(x,x^{\prime})=p_{x}(x-x^{\prime})/\hbar+p_{z}L/\hbar-\omega t,$ (21)
where $\omega=(p_{x}^{2}+p_{z}^{2})/(2m\hbar)$, and $t$ is the propagation time. If we use the classical expressions $p_{x}/p_{z}=(x-x^{\prime})/L$ and $t=mL/p_{z}$, then after subtracting off common phase contributions, we arrive at
$\Phi_{0}(x,x^{\prime})=\frac{p_{z}(x-x^{\prime})^{2}}{2L\hbar}.$ (22)
Since $p_{z}$ is fixed, there is a unique classical trajectory and phase for each pair of $x$ and $x^{\prime}$. The use of Eq. 22 can be easily justified for $p_{z}\gg p_{x}$ [38] or more generally by using path-integral theory [39].
The wavefunction $\psi(x)$ at some point on the detector is found by integrating the contributions from all source points $x^{\prime}$. If there is no photon scattered, the wavefunction shares a common phase at all points on the apertures and we have
$\psi(x)\propto\int_{\text{M}}\text{d}x^{\prime}e^{i\Phi_{0}(x,x^{\prime})}$ (23)
@@
$\psi_{n}(x)\propto\int_{\text{M}}\text{d}x^{\prime}\mbox{$\langle\gamma_{n}|\gamma_{x^{\prime}}\rangle$}e^{i\Phi_{0}(x,x^{\prime})},$ (24)
where $|\gamma_{x^{\prime}}\rangle$ is a photon emitted from $x^{\prime}$. We use the limit $\lambda\gg a$, so we treat the integral as a sum over the two point-like apertures.
Expressions for the matrix elements to be used in Eq. 24 are derived in Ref. [13]. We calculate atom patterns for the case of a circularly polarized incoming photon. For atom wavefunctions correlated with photon imaging-configuration position $X$ we use their Eq. 49 to find $\langle\gamma_{\text{X}}|\gamma_{x^{\prime}}\rangle$. For atom wavefunctions correlated with scattered photon direction, we have
$\mbox{$\langle\gamma_{\vec{k}_{\text{f}}}|\gamma_{x^{\prime}}\rangle$}=S(\vec{k},\vec{\kappa})e^{i\kappa_{x}x^{\prime}},$ (25)
where $S$ is the scattering amplitude.
@@
$\mathcal{P}_{x,\kappa_{x}}\propto(2k^{2}+\kappa_{x}^{2}-2k\kappa_{x})\left|\int_{\text{M}}\text{d}x^{\prime}e^{i\kappa_{x}x^{\prime}}e^{i\Phi_{0}(x,x^{\prime})}\right|^{2}.$ (26)
Again, in our limit $\lambda\gg a$, the integral can be treated as a sum over point-like slits, and we do so for simplicity. We numerically integrating over $k_{\text{f}x}$ to find the atom distribution $\mathcal{P}_{x}$.
Finally, far-field expressions used in App. B follow from the above expressions. For a photon with definite $\vec{k}_{\text{f}}$, we use Eq. 25 in Eq. 24. If $L\gg d$, then $p_{x}\approx(p_{z})(x/L)$, and the phase difference between the two sources is given by
$\phi^{\text{ff}}=p_{x}d/\hbar-\kappa_{x}d.$ (27)
@@
$\phi_{0}^{\text{ff}}=p_{x}d/\hbar,$ (28)
and
$\phi_{0}^{\text{ff}}=\frac{p_{z}d}{L\hbar}x.$ (29)
These are the well-known expressions for the far-field phase of an unperturbed two-slit interferometer.
Appendix B Semiclassical Pitfalls and Proper Calculation of Recoil Phase Change
First, we outline a sometimes-used sloppy argument, which arrives at the wrong result despite initial appearances of success. Consider a detector in the far-field limit. Since recoil takes $p_{x}\rightarrow p_{x}+\hbar\kappa_{x}$, we might be tempted to modify the well-known Eq. 28 by writing $\phi_{0}^{\text{ff}}\rightarrow\phi_{0}^{\text{ff}}+\kappa_{x}d$. From this expression, we could mistakenly conclude that we have arrived at the anticipated result, that randomizing recoil angles washes out atom fringes for $\lambda<d$. But it is the final momentum rather than the initial momentum which determines where the atom lands on the detector. What we actually calculated with the above transformations was simply how the original unperturbed fringe pattern varies in phase as we vary $p_{x}$, and with it the corresponding detector position $x$. There is no predicted fringe washout.
Continuing to make the mistake of neglecting recoil phase, we can be less sloppy than above by using Eq. 29. We might wonder whether the longitudinal component of recoil achieves the expected result. The substitution $p_{z}\rightarrow p_{z}+\hbar\kappa_{z}$ predicts a phase smearing $\Delta\phi=(x/L)(\kappa_{z}d)$. This can be understood as transit time broadening destroying the one-to-one correspondence between $p_{x}$ and $x$. This longitudinal decoherence is real, but it is not important for $x/L<\lambda/d$. If it were the whole story, there would still be of order $p_{z}/(\hbar k)\gg 1$ high-contrast fringes at small atom angles, even for $\lambda<d$.
Finally, we properly include recoil phase to find how the post-recoil and pre-recoil interferometer phases are related. For an atom with pre-recoil transverse momentum $p_{x_{0}}$, the far-field pre-recoil interferometer phase is $\phi_{0}^{\text{ff}}=p_{x_{0}}d/\hbar$. As discussed in Sec. III, recoil introduces a phase difference between atom wavefunctions leaving the two slits, and the post-recoil interferometer phase is given in Eq. 27. Recoil deflects the atom to a new angle, and using $p_{x}=p_{x_{0}}+\hbar\kappa_{x}$, we find $\phi^{\text{ff}}=\phi_{0}^{\text{ff}}$ because the $\kappa_{x}$ terms cancel. Thus for a two-slit interferometer, a deflected atom keeps its pre-recoil far-field interferometer phase, as Feynman implied. This result predicts fringe washout for $\lambda<d$, whereas the misguided approaches neglecting recoil phase do not.

### v3

@@ XII Conclusion
Decoherence is the penalty we pay for trying to separate out one part of a non-separable entangled state. Wonderfully, this “illegal” separation accomplishes what Copenhagen clumsily attempted by imposing its shifty split. The quantum freedom to choose different bases creates an ambiguity as to the facts of any unfolding story. In Feynman’s light microscope, is the atom following a classical single-path trajectory or a non-classical two-path trajectory? With full environmental information, both narratives are legitimate and both make experimentally verifiable predictions. But without ever imposing a microscopic/macroscopic distinction, decoherence causes plain quantum mechanics to pick out the classical single-path storylines as the only ones which can be recorded and used to make predictions in a complex environment. These decohered storylines each contain a set of unambiguous and causally connected facts.
@@
Alternatively, epistemic interpretations understand the wavefunction to be only a tool for making predictions. If the wavefunction never described microscopic reality, there is no basis for saying that any of the decohered branches predicted by unitary evolution actually exist, other than the one we are experiencing.
As a practical matter, MWI and epistemic approaches use identical equations, naturally include decoherence in identical ways, and make identical predictions for experiments. Opinions differ on whether the distinction is therefore an unimportant matter of philosophical preference [33, 39]. However, MWI and epistemic theories are together testable. For instance, if interference fringes of a big object were found to wash out for an inexplicable reason, then their underpinning unitary evolution would be falsified.
It is of course possible that quantum mechanics, and its extension quantum field theory, need to be modified in some fundamental way. Advocates point out that quantum theory has not yet been reconciled with gravity, and that there are still orders of magnitude between where it has been tested and the scale of macroscopic detectors. Alternate theories deserve to be tested when possible, but we should keep in mind that unitary evolution of Schrödinger’s equation has continued to prove successful under increasingly stringent conditions [2, 3, 4, 5]. Quantum mechanics has no crisis which proposed modifications would address, other than our human crisis of being forced to question cherished notions of reality or science.
Half a century ago, physicists could believe that wavefunctions described the reality of small things and that there was only one macroscopic reality. But after the development of decoherence theory and experimental scrutiny at relatively large mass scales, we can no longer rely on wavefunction collapse to provide that comfortable scenario. We must either say that plain quantum mechanics correctly predicts detector clicks but makes no statements about microscopic reality, or that it predicts a vast number of macroscopic realities besides the one we are currently experiencing.
@@ Acknowledgements.
We gratefully acknowledge stimulating and clarifying discussions with Isaac Cross, Jerry Gabrielse, Tim Kovachy, Jay Lawrence, and Bill Wootters.
@@ Appendix A Calculation of Interference Patterns
We are interested in obtaining atom patterns for a detector closer than the far-field limit, such that the single-slit patterns are somewhat separated from one another. To accommodate this requirement, we use a path integral approach. The phase acquired by an atom momentum eigenstate traveling from a source point $(x^{\prime},y^{\prime},0)$ to a detector point $(x,0,L)$ is given by
$\Phi(x,x^{\prime},y^{\prime})=\frac{p_{x}}{\hbar}(x-x^{\prime})-\frac{p_{y}}{\hbar}y^{\prime}+\frac{p_{z}}{\hbar}L-\omega t,$ (21)
where $\omega=(p_{x}^{2}+p_{y}^{2}+p_{z}^{2})/(2m\hbar)$, and $t$ is the propagation time. (Note that the presence of recoil has turned the problem, for the moment, from a 2D one to a 3D one, and a range of $y^{\prime}$ sources can contribute to detection at $y=0$.) If we use the classical expressions $p_{x}/p_{z}=(x-x^{\prime})/L$, $p_{y}/p_{z}=-y^{\prime}/L$, and $t=mL/p_{z}$, then after subtracting off common phase contributions, we arrive at
$\Phi(x,x^{\prime},y^{\prime})$ $=\frac{p_{z}}{2\hbar L}\left((x-x^{\prime})^{2}+(y^{\prime})^{2}\right)$ (22)
$=\frac{p_{0}+\hbar k_{z}}{2\hbar L}\left((x-x^{\prime})^{2}+(y^{\prime})^{2}\right).$
The $k_{z}$ term is of order $\hbar^{2}kL/(p^{2}_{0}a^{2})$ and the $y^{\prime}$ term is of order $\hbar k^{2}L/p_{0}$. By increasing $p_{0}/(\hbar k)$, these terms can be made arbitrarily small. In other words, for sufficiently short transit times, the photon scattering perturbation has negligible effect on the phase accumulated over transit. We use this high-momentum limit to simplify the calculations, and we have
$\Phi(x,x^{\prime})=\frac{p_{0}}{2\hbar L}(x-x^{\prime})^{2}.$ (23)
The wavefunction $\psi(x)$ at some point on the detector is found by integrating the contributions from all source points $x^{\prime}$. For this purpose, the use of classical trajectories, used to derive Eq. 23, can be easily justified for $p_{0}\gg p_{x}$ [40] or more generally by using path-integral theory [41]. If there is no photon scattered, the wavefunction shares a common phase at all points on the apertures, and we have
$\psi(x)\propto\int_{\text{M}}\text{d}x^{\prime}e^{i\Phi(x,x^{\prime})},$ (24)
@@
$\psi_{n}(x)\propto\int_{\text{M}}\text{d}x^{\prime}\mbox{$\langle\gamma_{n}|\gamma_{x^{\prime}}\rangle$}e^{i\Phi(x,x^{\prime})},$ (25)
where $|\gamma_{x^{\prime}}\rangle$ is a photon emitted from $x^{\prime}$.
Expressions for the matrix elements to be used in Eq. 25 are derived in Ref. [13]. For atom wavefunctions correlated with photon imaging-configuration detection at $x_{\gamma}$, we use their Eq. 49 to find $\langle\gamma_{x_{\gamma}}|\gamma_{x^{\prime}}\rangle$. For atom wavefunctions correlated with scattered photon direction, we have
$\mbox{$\langle\gamma_{\vec{k}_{\text{f}}}|\gamma_{x^{\prime}}\rangle$}=S(\vec{k}_{\text{f}})e^{i\kappa_{x}x^{\prime}},$ (26)
where $S$ is the scattering amplitude for a photon scattered from $x=0$.
@@
$\mathcal{P}_{x,\kappa_{x}}\propto(2k^{2}+\kappa_{x}^{2}-2k\kappa_{x})\left|\int_{\text{M}}\text{d}x^{\prime}e^{i\kappa_{x}x^{\prime}}e^{i\Phi(x,x^{\prime})}\right|^{2}.$ (27)
This equation corresponds to Eq. 53 of [13], whose procedure we followed, but we have generalized their expression to accommodate our finite aperture width and non-far-field atom detection. We numerically integrating over $k_{\text{f}x}$ to find the atom distribution $\mathcal{P}_{x}$. Since we only consider $\lambda\gg a$, we simplify the computation by treating the recoil term as constant over each slit.
Finally, narrow-slit expressions for interferometer phase, used in App. B, follow from the above expressions. For a photon with definite $\vec{k}_{\text{f}}$, we use Eq. 26 in Eq. 25. The phase difference between narrow slits is $\Delta\phi=\phi_{\text{L}}-\phi_{\text{R}}$, with the sign chosen to make the interferometer phase increase as $x$ increases. Using $\phi_{\text{L,R}}=(p_{0})/(2\hbar L)(x\pm d/2)^{2}\mp\kappa_{x}d/2$, we get
$\Delta\phi$ $=\frac{p_{0}d}{L\hbar}x-\kappa_{x}d$ (28)
$=\frac{p_{x}d}{\hbar}-\kappa_{x}d.$ (29)
@@
$\Delta\phi_{0}$ $=\frac{p_{0}d}{L\hbar}x$ (30)
$=\frac{p_{x}d}{\hbar}.$ (31)
These are the well-known expressions for the phase of an unperturbed two-narrow-slit interferometer.
Appendix B Semiclassical Pitfalls and Proper Calculation of Deflection Phase Change
First, we outline a sometimes-used sloppy argument, which arrives at the wrong result despite initial appearances of success. Since recoil takes $p_{x}\rightarrow p_{x}+\hbar\kappa_{x}$, we might be tempted to modify the well-known Eq. 31 by writing $\Delta\phi_{0}\rightarrow\Delta\phi_{0}+\kappa_{x}d$. From this expression, we could mistakenly conclude that we have arrived at the anticipated result, that randomizing recoil angles washes out atom fringes for $\lambda<d$. But it is the final momentum rather than the initial momentum which determines where the atom lands on the detector, and the above argument forgets that as we change $p_{x}$ we also change the detection position $x$. What we actually calculated was simply how the original unperturbed fringe phase changes with $x$ and its corresponding $p_{x}$. There is no predicted fringe washout.
Continuing to make the mistake of neglecting recoil phase, we can be less sloppy than above by using Eq. 30. We might wonder whether the longitudinal component of recoil achieves the expected result. The substitution $p_{0}\rightarrow p_{0}+\hbar\kappa_{z}$ predicts a phase smearing $\delta\phi=(x/L)(\kappa_{z}d)$. This can be understood as transit-time broadening destroying the one-to-one correspondence between $p_{x}$ and $x$. This type of longitudinal decoherence is real, but it is not important for $x/L<\lambda/d$. If it were the whole story, there would still be of order $p_{0}/(\hbar k)\gg 1$ high-contrast fringes at small atom angles, even for $\lambda<d$.
Finally, we properly include recoil phase to find how the post-recoil and pre-recoil interferometer phases are related. For an atom with pre-recoil transverse momentum $p_{x_{0}}$, the pre-recoil interferometer phase (Eq. 31) is $\Delta\phi_{0}=p_{x_{0}}d/\hbar$. As discussed in Sec. III, recoil introduces a phase difference between atom wavefunctions leaving the two slits, and the post-recoil interferometer phase is given by Eq. 29. Recoil deflects the atom to a new angle, and using $p_{x}=p_{x_{0}}+\hbar\kappa_{x}$, we find $\Delta\phi=\Delta\phi_{0}$ because the $\kappa_{x}$ terms cancel. Thus for a two-slit interferometer, a deflected atom keeps its pre-recoil interferometer phase, as Feynman implied. This result predicts fringe washout for $\lambda<d$, whereas the misguided approaches neglecting recoil phase do not.